OpenStudy (anonymous):

If y varies inversely as x, find the constant of variation if y = 14 when x = 8.

OpenStudy (johnweldon1993):

\[\large y = \frac{k}{x}\] When y = 14, x = 8 so plug those in \[\large 14 = \frac{k}{8}\] \[\large k = 14 \times 8 = \space ?\]

OpenStudy (anonymous):

112?

OpenStudy (johnweldon1993):

Looks right to me!